0 0
Add a comment Improve this question Transcribed image text
Answer #1

The momnet of inertia of the hoop is,

$$ I_{h o 0 p}=m R^{2} $$

The momnet of inertia of the bowling ball is,

$$ I_{\text {sowing } \operatorname{gal}}=\frac{2}{5} m R^{2} $$

The momnet of inertia of the penny is,

$$ I_{\text {yenyy }}=\frac{1}{2} m R^{2} $$

The momnet of inertia of the hollow globe is,

$$ I_{\text {houlowglade }}=\frac{2}{3} m R^{2} $$

The acceleration of the object when it is moving from inclined path is indirectly proportional to the mom ent of inertia. If all the objects are having same mass and moving with the same linear speed, then from the above calculations, the rolling speed is order from fasstest to slowest is.

bowling ball \(>\) penny \(>\) hollow globe \(>\) hula hoop